End Of Game What is the SSE? (numerical answer - round to one decimal place) NEXT BACK G Search or type URL
End Of Game Anova: Two-Factor Without Replication SUMMARY Count Sum Average Variance 1 5 11322.6 22.3 | 23 | 10.51 5 115 3 5 12224.412.8 5 127254 21.8 25.4 132 26.4 32.3 64.3 5 128 5110 25.6 Mon Tue Wed Thurs Fri 7 175 7 168 25 40 24 17 7 14721 3.333333 22 6.666667 7203 29 13.33333 71 154 ANOVA Source of Variation MSF P-value F crit 0.863774 0.535234 2.508189 Rows Columns Error 85.6 271.6 67.9 24 Total 753.634 DONE
End Of Game What is the SSE? (numerical answer - round to one decimal place) NEXT BACK G Search or type URL
iti tml # / results MobLab Student Con End Of Game 2 What is the MSBL? (numerical answer -round to two decimal places) Input Answer NEXT BACK
My ASU results x MobLab dent Console End Of Game What is the MSE? (numerical answer - round to two decimal places) Answer DONE K BACK G Search or type URL
Answer #1

Solution:

First, we calculate the degrees of freedom (df) of columns

MS (columns) = SS (columns)/df (columns)

67.9 = 271.6/df (columns)

df (columns) = 4

Now, we find the df (rows)

df (total) = df (rows) + df (columns) + df (error)

34 = df (rows) + 4 + 24

df (rows) = 6

MS (rows) = SS (rows)/df (rows)

MS (rows) = 85.6/6

MS (Rows) = 14.267

SS (Error) = SS (total) - SS (Rows) - SS (Columns)

SS (Error) = 753.6 - 85.6 - 271.6

SS (Error) = 396.4

MS (Error) = SS (Error)/df (Error)

MS (Error) = 396.4/24

MS (Error) = 16.52